# Why Johnny Signs with Sigstore: Examining Tooling as a Factor in Software Signing Adoption in the Sigstore Ecosystem

### v3

@@ 6 Discussion
Privacy Concerns: Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern. This issue also affects the use of OIDC for keyless signing. While this challenge is not unique to Sigstore [52] and is a recurring issue in software signing and modern cryptography, it underscores the need for continued innovation to address privacy-preserving mechanisms that balance transparency with user privacy, ensuring broader adoption and trust in cryptographic tools like Sigstore. While works like Speranza[52] has been proposed, efforts should focus on integrating these solutions to tools like Sigstore.
7 Conclusion
Software signing allows software provenance to be guaranteed through cryptographic primitives. This property would address many software supply chain attacks. However, for decades, the widespread adoption of software signing has eluded the software industry. In this study, we examined a recent change in software signing practices by focusing on Sigstore, the leading next-generation software signing platform. We interviewed industry experts to understand Sigstore from a usability perspective. We identified the factors they weigh and the challenges they face. Our results shed light on usability issues for software signing in particular, and will inform more usable software tools writ broad.
8 Acknowledgments
@@
Appendix A Interview Protocol
Table 1 gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
We include Section D of the protocol for completeness, although it was not analyzed in this study.
Table 7: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions. We only analyze parts A and D in this study
Themes Sub-themes Questions
A: Demographic A-1: What best describes your role in your team? (Security engineer, Infrastructure, software engineer, etc)
A-2: What is your seniority level? How many years of experience?
A-3: What is the team size?
A-4: $\dagger$ What are the team’s major software products/artifacts?
A-5: What type of Organization/company?
@@ Appendix A Interview Protocol
Appendix B Codebooks
Table 8 describes our codebook, with code definition, and example quote tagged with that code.
Table 8: Excerpts from the final codebook we used for our analysis.
@@ Appendix B Codebooks
Sigstore Criticisms Criticisms and weaknesses associated with implementing Sigstore for signing. S3: "We do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. "
@@
Sigstore Improvements Suggested improvements for Sigstore. S9: "I would say probably integrations within different platforms, so say you use GitLab or GitHub or Jenkins or something like that. Try to leverage some type of integration to make that, I guess, less friction for the developers to implement, something like that."
@@
Sigstore Strengths Strengths of Sigstore and reasons why subjects prefer using Sigstore. S7: "The convenience and the good user experience is probably what keeps us using Sigstore "
@@
Sigstore Alternatives Considered Alternatives of Sigstore that were or are being considered by the subject (for subjects whose primary signing implementation is Sigstore). S5: "We are also looking at other new alternatives that are coming up, like OpenPubkey. Docker just released a new standard called OpenPubKey, which is trying to solve this issue where they will not require you to have infrastructures and stuff."
@@
Sigstore Components Used Components of Sigstore (Cosign, Fulcio, etc.) being used by the subject. S2: "Mostly fulcio is what’s integrated into our product."
@@
Notary Criticism Criticisms and weaknesses Associated with Notary – promoting users to stop using them. "S5: "We were using Notary, and Notary only verifies that the image was signed using a specific key, but you cannot verify that who is the owner of the key."
Appendix C Traditional vs Next-Gen Software Signing Workflow
We show a typical traditional package signing workflow in Figure 4 compared to a next-generation workflow in Figure 5.
Figure 4: Typical workflow for software signing and verifying signatures. The component author packages (A) and signs (B) their software. The signed package (C) and public key (D) are published. To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
Figure 5: Typical workflow for next-generation software signing. The component author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider. The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
Traditional Signing: The signature creation process begins when the maintainer’s artifact is ready for submission. The signer generates a key pair that provides them with a private key for signing their software artifact and a public key for a verifier to validate the artifact’s signature. Once the artifact is signed, it must be submitted to a package registry (such as PyPi, npm, or Maven) or the intended party. During this phase, the signer is responsible for making their public key available and discoverable so future users can verify the signer’s identity. In the verification phase, the verifier retrieves the signer’s public key, uses it to decrypt the signature file, and checks for equivalency with the software artifact’s hashed digest.
Next-generation Signing:: Keyless/Ephemeral software signing mitigates the security risk of private key compromise. Keyless software signing differs from traditional public-key methodology in the fact that the signer does not rely on a long-term private key to sign their messages. Instead of the signer being responsible for the secure maintenance of a private key, in keyless signing, the signer uses a short-lived key that only lasts long enough to sign their artifact [2].
Appendix D Survey Attempt
@@
Appendix E Code Saturation
We present a saturation analysis of our studies in  Figure 6. The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview. We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 11. Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4. This is also reflected in the orange trendline (unique number of codes).
Appendix F Code Evolution Diagram
Figure 7 describes the evolution of our codebook over six rounds of revisions.
Figure 7: Codebook Evolution: The codebook underwent six rounds of revisions, with changes highlighted by different colors: round 2 (pink), round 3 (blue), round 4 (green), round 5 (yellow), and round 6 (purple). The final codes from round 6 were used in this paper, and categorized thematically. We combined framework and thematic analyses to develop the codebook. First, we induced codes from each transcript and analyzed them thematically. During the framework analysis, we mapped codes to the four stages/blocks of the software supply chain factory model. Round 1: The codebook was defined based on emerging themes. Round 2: No changes. Round 3: Blue-highlighted codes (5) were added after memoing and coding the second set of transcripts. Round 4: Green-highlighted codes (2) were added. Round 5: No changes. Round 6: Analysts A1 and A2 determined some codes should be further divided. Purple-highlighted codes (4) were added and the codebook was finalized. We focused on the topics of relating to sigstore(signing tools) including, Notary, Vault, Spiffe, GPG, and others.

### v4

@@ VI-C Privacy Concerns
Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern. Privacy concerns were also major issues influencing users not to consider sigstore as suitable for enterprise. This issue also affects the use of OIDC for keyless signing. While this challenge is not unique to Sigstore [71] and is a recurring issue in software signing and modern cryptography, it underscores the need for continued innovation to address privacy-preserving mechanisms that balance transparency with user privacy, ensuring broader adoption and trust in cryptographic tools like Sigstore. While works like Speranza[71] has been proposed, efforts should focus on integrating these solutions to tools like Sigstore.
VII Threats to Validity
To discuss the limitations of our work, we follow Verdecchia et al.’s [90] recommendation to reflect on the greatest threats to our research. In addition, we discuss how our experiences and perspectives may have influenced our research (positionality).
Construct Validity: Cresswell’s framework targets health IT, where “tooling” denotes integrated clinical systems, so applying it to software signing carries construct risks. We mitigated risks by inductively validating themes (from its core factors) against practitioner interviews, ensuring our adapted constructs reflect signing-tool usability.
External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore. We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature. A second threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects. We attempted a survey (details are in section -E) but received only a few responses, echoing the challenges of surveying specialized communities [91]. A third threat is our study’s temporal scope. Interviews were conducted between November 1, 2023, and February 9, 2024, capturing Sigstore users’ experiences during that period. Since then, Sigstore has introduced some changes, but not the core workflows and components. This suggests our findings remain relevant.
Statement of Positionality: The author team has expertise in software supply chain security and qualitative research methods, e.g., publishing top-tier papers in software engineering and cybersecurity venues. One of the authors is a Sigstore contributor. To avoid bias, this author was not involved in developing the research design nor the initial analysis of the results, but provided insight in later analysis.
VIII Conclusion
Software signing guarantees provenance via cryptographic primitives, reducing supply-chain attacks. Next-generation tools like Sigstore automate traditional challenges, yet their usability remains underexplored. We interviewed practitioners to examine Sigstore’s usability, identifying key adoption factors. Our findings show that Sigstore’s identity-based workflows improve usability, but integration complexities, privacy concerns, and organizational constraints hinder adoption. We recommend toolmakers provide official integration plugins and privacy controls for transparency logs. These insights inform automation improvements, promoting secure, usable next-generation signing in automated software engineering.
IX Acknowledgments
@@
-C Codebooks
Table VII describes our codebook, with code definition, and example quote tagged with that code.
TABLE VII: Excerpts from the final codebook we used for our analysis.
@@
Sigstore Criticisms Criticisms and weaknesses associated with implementing Sigstore for signing. P3: "We do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. "
@@
Sigstore Improvements Suggested improvements for Sigstore. P9: "I would say probably integrations within different platforms, so say you use GitLab or GitHub or Jenkins or something like that. Try to leverage some type of integration to make that, I guess, less friction for the developers to implement, something like that."
@@
Sigstore Strengths Strengths of Sigstore and reasons why subjects prefer using Sigstore. P7: "The convenience and the good user experience is probably what keeps us using Sigstore "
@@
Sigstore Alternatives Considered Alternatives of Sigstore that were or are being considered by the subject (for subjects whose primary signing implementation is Sigstore). P5: "We are also looking at other new alternatives that are coming up, like OpenPubkey. Docker just released a new standard called OpenPubKey, which is trying to solve this issue where they will not require you to have infrastructures and stuff."
@@
Sigstore Components Used Components of Sigstore (Cosign, Fulcio, etc.) being used by the subject. P2: "Mostly fulcio is what’s integrated into our product."
@@
Notary Criticism Criticisms and weaknesses Associated with Notary – promoting users to stop using them. "P5: "We were using Notary, and Notary only verifies that the image was signed using a specific key, but you cannot verify that who is the owner of the key."
-D Code Saturation
We present a saturation analysis of our studies in  Figure 7. The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview. We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 12. Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4. This is also reflected in the orange trend line (unique number of codes).
-E Survey Attempt
@@
-F Sigstore Components Most Commonly Used by Our Sample Population
TABLE VIII: Sigstore Functionalities Used by Subjects.
Tool # of Subjects
Cosign (Signing & Verification CLI) 9
Certificate Authority (Fulcio) 9
Keyless Signing & Identity Manager (OIDC) 9
Transparency Log (Rekor) 8
Gitsign (Commit Signing) 4
Component Customization & Local Sigstore Deployment 2
-G Expanded Demographics Table
TABLE IX: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff. We highlight the type of software subject’s immediate team produces. * refers to cases where subject’s team built other types of software in addition to the main stated. O refers to cases where subject belonged to multiple teams with other products than the one stated.
ID Role Experience Software Type
P1 Research leader 5 years Internal POC software
P2 Senior mgmt. 15 years SAAS security tool*O
P3 Senior mgmt. 13 years SAAS security tool*O
P4 Technical leader 20 years Open-source tooling
P5 Engineer 2 years Internal security tooling
P6 Technical leader 27 years Internal security tooling*O
P7 Manager 6 years Security tooling*O
P8 Technical leader 8 years Internal security tooling*
P9 Engineer 2.5 years SAAS security*
P10 Engineer 13 years SAAS security*
P11 Technical leader 16 years Firmware*
P12 Technical leader 4 years Consultancy
S13 Senior mgmt. 16 years Internal security tool
P14 Research leader 13 years POC security Software*
P15 Senior mgmt. 15 years Internal security tooling
P16 Senior mgmt. 15 years SAAS
P17 Manager 11 years Security tooling
-H Factors Influencing Sigstore Adoption (Sigstore Users)
TABLE X: Reasons Practitioners Choose or Switch to Sigstore Before Adoption. Practitioners’ decisions are driven primarily by human/social factors—such as community contribution, frustration with legacy tools, and usability issues—supplemented by Sigstore’s technological capabilities and reinforced by macroenvironmental pressures like regulations and trust in the CNCF ecosystem.
Topics & Associated Examples Subjects
Human/Social Factors
Practitioners Contribute to Sigstore 6 – P2, P4, P6, P7, P9, P14
GPG Issues 6
Low adoption rates P1, P14
Key management issues & Other usability concerns P1, P6, P12, P16, P17
Steep learning curve P12, P16
Compatibility with newer technology P16
Notary Issues 2
Non-demand from customers P12
Compatibility with other tools P12
Lack of regular updates P12
Key & Identity Management P5
Proprietary Tool Issues 1
Difficult to setup P15
Technological Factors
Available Sigstore Functionalities 3 – P1, P5, P14
A transparency log, etc P5, P14
Compatibility to other Tools P1
Macroenvironmental Factors
Regulation & Standards 4 – P5, P6, P15, P17
Large User Community 1 – P9
Inherent Trust of Creators 1
Trust of CNCF products P3
-I Factors Influencing Sigstore Consideration (Non-Sigstore Users).
TABLE XI: Non Sigstore Users: Tooling Issues causing a consideration of Sigstore and other Next-gen tools.
Tool Subjects Topic
Internal tool P8, P11 Tooling Issues
(1) Integration with other tools/platformsT – P8, P11
(2) Non unified tool implementation across teams O – P8, P11
(3) Signer ID Management T – P11
(4) Usage info/documentation unclear T – P11
(5) Verification not built in with tool O – P8
(6) Non transparency in signing infrastructure T – P11
(7) Security of Signing Infrastructure T – P11
(8) Engineers don’t understand the current tool P – P11
Considered (or Adopted) Fix
(1) Organization building new tools with Next-gen features – P8, P11
Notary-V1 P10 Tooling Issues
(1) Key ManagementT
(2) Lack of an inbuilt signing policy T
(3) Engineers don’t understand the current tool H – P11
(4) Integration with other tools/platformsT
Considered (or Adopted) Fix
(1) Organization has adopted Openpubkey for key management
(1) Organization considering Sigstore, and Notary V2
PGP P13 (1) Tooling Issues
(1) Key ManagementT
(2) Non transparency in signing infrastructure T
(3) Security of Signing Infrastructure T
Considered (or Adopted) Fix
(1) Use of Key discovery and delivery systems like TUF [^4] The Update Framework[93]
-J Additional Discussions & Implications of Results
In addition to the implications discussed in section VI, we outline further implications of our findings below.
-J1 Generalizability of Findings to Other Classes of Tools in Software Engineering
Our findings reinforce several adoption factors noted in prior literature on open source tools—such as ease of use, compatibility with existing technologies, and integration flexibility—which were especially salient for next-generation signing tools like Sigstore. They also highlight additional considerations that may generalize to other classes of secure software engineering tools example the role played by macroenvironmental factors like the tooling community, and regulations.
Furthermore, our results suggest adoption patterns not widely discussed in existing research. In particular, we observe the emergence of a migration pathway model, where teams incrementally adopt next-generation tools. Several non-Sigstore users reported experimenting with or partially integrating smaller solutions—such as OpenPubKey and TUF—to meet specific needs like key delivery and management. This suggests a stepwise transition process in tooling adoption that may be common across emerging tool ecosystems. This model can also be important to understand and further propose a streamlined organizational adoption pathway for software engineering tools.
-J2 The Surprising Role of Community in Software Signing Tool Adoption
Many participants cited their involvement in the Sigstore community as a key reason for adopting the tool (Table X). While this may partly reflect our sampling, half of the non-Sigstore users who adopted external tools also emphasized community affiliation as a motivator. Challenges like setting up private Sigstore instances or accessing usage knowledge were frequently linked to the absence of sufficient community support (section V-A2). Other community-related influences, such as the size of the user community and trust in the tool creators, also shaped adoption decisions.
These findings echo prior work on the importance of community in developer tool adoption [94]. Toolmakers should actively foster supportive, trustworthy communities around their tools, as these networks can be pivotal to both adoption and long-term sustainability.
-K Legible Images
We have added a more legible version of our Sigstore workflow imageFigure 3 here Figure 8.
Figure 8: Sigstore Signing Workflow. The software author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider. The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.